### v1

@@ 3 Inexact Proximal Gradient Method
Recall that, given $x_{k}\in\mathbbm{E}$, the exact solution of the proximal gradient subproblem (3) with $\gamma_{k}=1$ consists of finding $\bar{x}_{k}$ such that
$\bar{x}_{k}={\rm arg}\!\min_{x\in\mathbbm{E}}\left\{\left\langle\nabla f(x_{k}),x-x_{k}\right\rangle+g(x)+\frac{1}{2}\|x-x_{k}\|^{2}\right\}.$
This proximal subproblem is equivalent to the following monotone inclusion problem:
$0\in\nabla f(x_{k})+\partial g(\bar{x}_{k})+\bar{x}_{k}-x_{k}.$ (9)
The $k$-th iteration of the inexact proximal gradient method, which incorporates an explicit linesearch as proposed below, begins by first relaxing the above inclusion. This is achieved by finding an approximate solution $\tilde{x}_{k}$ along with a residual pair $(v_{k},\varepsilon_{k})$ satisfying
$v_{k}\in\nabla f(x_{k})+\partial_{\varepsilon_{k}}g(\tilde{x}_{k})+\tilde{x}_{k}-x_{k},$
and adheres to a specific mixed relative error criterion. This criterion is designed to ensure that the residual pair remains small, while also controlling the functional value of $g$ and the angle between $\nabla f(x_{k})$ and $v_{k}$, in addition to guiding a search direction that involves the intermediate iteration $\tilde{x}_{k}$ and the residual $v_{k}$. If the stopping criterion $\tilde{x}_{k}=x_{k}$ is not satisfied, we use a novel linesearch procedure, designed with the triple $(\tilde{x}_{k},v_{k},\varepsilon_{k})$ in mind, to find a suitable stepsize $\beta_{k}$. This stepsize is used to move from the current point $x_{k}$ towards $y_{k}-x_{k}$ where $y_{k}=\tilde{x}_{k}-v_{k}$, genegenerating the next point $x_{k+1}$.
In the following, we formally describe the proposed method, which combines a relative inexact computation of the proximal subproblem with an explicit linesearch.
@@
Remark 1 (The Inexact Proximal Subproblem)
Note that the inexact proximal gradient subproblem (10) is equivalent to the exact proximal gradient subproblem (9) whenever $v_{k}=0$ and $\varepsilon_{k}=0$. Moreover, it is possible to always set $v_{k}$ to 00 and still introduce a novel scheme that diverges from the previous one, which was based on inexact summable error tolerance. Indeed, under this assumption, the inexact proximal subproblem (10)-(11) can be rewritten as:
$\varepsilon_{k}\leq\frac{(1-\tau-\alpha)}{2(1+\gamma_{2})}\|x_{k}-\tilde{x}_{k}\|^{2}.$ (14)
It is worth noting that for the above subproblem, no summability assumption is made on the tolerance sequence $(\varepsilon_{k})_{k\in\mathbbm{N}}$ to achieve the convergence and complexity of the proposed method, as will be shown later. Furthermore, inclusion (10) and inequality (11) guarantee that $\tilde{x}_{k}$ and $\tilde{x}_{k}-v_{k}$ are in the ${\bf dom\,}g$. Additionally, if $g$ is the indicator function of a nonempty closed convex set $C$, then (13)–(14) can be fulfilled by finding a point $\tilde{x}_{k}\in C$ such that
$\langle x_{k}-\nabla f(x_{k})-\tilde{x}_{k},y-\tilde{x}_{k}\rangle\leq\varepsilon_{k}:=\frac{(1-\tau-\alpha)}{2(1+\gamma_{2})}\|x_{k}-\tilde{x}_{k}\|^{2},\quad\forall y\in C.$ (15)
For example, if $C$ is bounded, one can use the conditional gradient (CondG) method, a.k.a. Frank-Wolfe method Frank:1956 ; Jaggi:2013 , to obtain such point $\tilde{x}_{k}\in C$. Indeed, given $z_{t}\in C$, the $t$-th step of the CondG method, applied to solve the projection problem $\min_{y\in C}\|y-x_{k}+\nabla f(x_{k})\|^{2}/2$, first finds $w_{t}$ as a minimum of the linear function $\langle{z_{t}-x_{k}+\nabla f(x_{k})},{\cdot-z_{t}}\rangle$ over $C$ and then sets $z_{t+1}=(1-\alpha_{t})z_{t}+\alpha_{t}w_{t}$ for some $\alpha_{t}\in(0,1)$. If at an iteration $\ell$, the computed point $z_{\ell}$ satisfies the stopping criterion $\langle{z_{\ell}-x_{k}+\nabla f(x_{k})},{w_{\ell}-z_{\ell}}\rangle\geq-\varepsilon_{k}$ for the CondG method, then $\tilde{x}_{k}:=z_{\ell}$ and (15) will hold.
@@
Note that the explicit linesearch proposed in Step 4 of the IPG-ELS is related to the one proposed in BelloCruz:2016a when $v_{k}=0$ and $\varepsilon_{k}=0$. Moreover, note that the linesearch does not evaluate the inexact proximal subproblem inside of the inner loop of Step 4.
@@
The subsequent analysis investigates the linesearch procedure introduced in Step 4 of the IPG-ELS method, assuming that the gradient of the function $f$, $\nabla f$, is Lipschitz continuous. We aim to establish an upper bound for the number of iterations required by the linesearch procedure in Step 4.
@@
This lemma provides a sufficient condition to ensure that the lower bound of the sequence generated by our linesearch is greater than 00. Specifically, if $\nabla f$ is $L$-Lipschitz, then the step sizes $\beta_{k}$, produced through the line search (12), are guaranteed to be bounded below by a positive constant $\beta>0$, i.e., $\beta_{k}\geq\beta>0$ for all $k\in\mathbb{N}$. Moreover, it is possible to relax the global Lipschitz condition to something local, such as $\nabla f$ being locally Lipschitz continuous around any solution of problem (1), as was done in Proposition 5.4(ii) of BelloCruz:2016a .
4 Convergence and Complexity Analyses of the IPG-ELS Method
In this section, we focus on analyzing the convergence properties of the IPG-ELS method. We establish the convergence and its iteration complexity in terms of functional values and a residual associated with an approximated solution, as defined in Definition 1.
@@ 4 Convergence and Complexity Analyses of the IPG-ELS Method
Additionally, the functional value sequence $(F(x_{k}))_{k\in\mathbbm{N}}$ is decreasing and convergent, and $\sum_{k=0}^{\infty}\|x^{k}-x^{k+1}\|^{2}<+\infty$.
@@
So, $(F(x_{k}))_{k\in\mathbbm{N}}$ is decreasing and convergent because it is bounded from below by $F(x_{*})$. Moreover, the last inequality implies that $\sum_{k=0}^{\infty}\|x^{k}-x^{k+1}\|^{2}<+\infty$. ∎
Next, we establish the convergence rate of the functional values sequence $(F(x_{k}))_{k\in\mathbbm{N}}$.
Theorem 4.1 (Convergence Rate of the IPG-ELS Method)
Summing the above inequality over $\ell=0,1,\ldots,k$, we have
$(k+1)\left(F(x_{k})-F(x_{*})\right)\leq\frac{1}{2\beta}\left(\|x_{0}-x_{*}\|^{2}+2\left[F(x_{0})-F(x_{*})\right]\right).$
Since $x_{*}\in S_{*}$ is arbitrary, the proof of (24) follows. ∎
Remark 3 (Complexity of $\eta$-Approximate Solution)
It follows from Theorem 4.1 that every accumulation point of the sequence $(x_{k})_{k\in\mathbbm{N}}$ is a solution to problem (1). Moreover, given any $\eta>0$, in at most $k=\mathcal{O}(1/\eta^{2})$ iterations, the IPG-ELS method generates an $\eta$-approximate solution $x_{k}$ to problem (1), in the sense that $F(x_{k})-F_{*}\leq\eta$.
Theorem 4.2 (Complexity of $\eta$-Approximate Stationary Solution)
Additionally, if $\alpha\in(0,1-\tau)$, and $\nabla f$ is Lipschitz continuous on ${\bf dom\,}g$, then, given a tolerance $\eta>0$, the IPG-ELS method generates an $\eta$-approximate stationary solution $\tilde{x}_{k}$ to problem (1) with residues $(w_{k},\varepsilon_{k})$, in the sense of Definition 1, in at most $k=O\left(1/\eta^{2}\right)$ iterations.
$\|\tilde{x}_{{\ell_{k}}}-x_{{\ell_{k}}}\|^{2}\leq\frac{d_{0}^{2}+2[F(x_{0})-F(x_{*})]}{\alpha\beta(k+1)}.$ (28)
Now, from the definition of $w_{\ell}$, (29), the Cauchy-Schwarz inequality and the Lipschitz continuity of $\nabla f$, we have, for every $\ell\in\mathbbm{N}$,
Moreover, it follows from (29) that
which in turn implies that
We end this section by establishing the full convergence of the sequence $(x_{k})_{k\in\mathbbm{N}}$ to a solution of problem (1). The proof is based on the quasi-Fejér convergence of the sequence $(x_{k})_{k\in\mathbbm{N}}$ to the set $S_{*}$, as defined in Definition 2.
@@
$\sum_{k=0}^{\infty}\delta_{k}=$ $2\sum_{k=0}^{\infty}\Big{[}F(x_{k})-F(x_{k+1})\Big{]}\leq 2\Big{[}F(x_{0})-\lim_{k\to\infty}F(x_{k+1})\Big{]}$
@@
This together with (33) tells us that the sequence $(x_{k})_{k\in\mathbbm{N}}$ is quasi-Fejér convergent to $S_{*}$ via Definition 2. By Lemma 1(a), the sequence $(x_{k})_{k\in\mathbbm{N}}$ is bounded. Let $\bar{x}$ be a cluster point of $(x_{k})_{k\in\mathbbm{N}}$. Hence there exists a subsequence $(x_{\ell_{k}})_{k\in\mathbbm{N}}$ of $(x_{k})_{k\in\mathbbm{N}}$ converging to $\bar{x}$.
@@
Summing for $k=0,1,\ldots,m$, the above inequality implies that
@@
$\beta\sum_{k=0}^{\infty}\left[F(x_{\ell_{k}})-F(x_{*})\right]\leq\sum_{k=0}^{\infty}\beta_{\ell_{k}}\left[F(x_{\ell_{k}})-F(x_{*})\right]\leq\sum_{k=0}^{\infty}\beta_{k}\left[F(x_{k})-F(x_{*})\right]<+\infty,$
@@
$F(x_{*})\leq F(\bar{x})\leq\liminf_{k\rightarrow\infty}F(x_{\ell_{k}})=\lim_{k\rightarrow\infty}F(x_{\ell_{k}})=F(x_{*}),$
@@
therefore, $\lim_{k\rightarrow+\infty}\,\|x_{\ell_{k}}-\tilde{x}_{\ell_{k}}\|=0$ and $\lim_{k\rightarrow+\infty}\|v_{\ell_{k}}\|=0$. Additionally, we can use (11) to show that $\lim_{k\rightarrow+\infty}\varepsilon_{\ell_{k}}=0$. Thus, $\bar{x}$ is also an accumulation point of the sequence $(\tilde{x}_{k})_{k\in\mathbb{N}}$, and without loss of generality, we can assume that $\tilde{x}_{\ell_{k}}\to\bar{x}$ as $k\to+\infty$. Moreover, we have that
@@ 5 Numerical Experiments
In this section, we investigate the numerical behavior of the IPG-ELS method in solving the CUR-like factorization optimization problem JMLR:v12:mairal11a . Consider $\mathbbm{E}=\mathbbm{R}^{n\times m}$. Given a matrix $W\in\mathbbm{R}^{m\times n}$, the objective is to find a sparse rows and columns matrix, $X\in\mathbbm{R}^{n\times m}$, such that $WXW$ approximates $W$. This problem can be formulated as the following splitting optimization problem:
$\min_{X\in\mathbbm{R}^{n\times m}}\frac{1}{2}\|W-WXW\|_{F}^{2}+\lambda_{\text{row}}\sum_{i=1}^{n}\|X^{(i)}\|_{2}+\lambda_{\text{col}}\sum_{j=1}^{m}\|X_{(j)}\|_{2},$
@@
In this case, the gradient of $f$ is given by $\nabla f(X)=W^{T}(W-WXW)W^{T}$ and has a Lipschitz constant $L=\|W^{T}W\|^{2}$. However, the proximal operator of $g$ does not have a closed-form solution. Following the approach in Schmidt:2011 (see also Barre:2022 ; Zhou:2022 ), a triple $(\tilde{X}_{k},V_{k},\varepsilon_{k})$ satisfying (10)-(11) can be obtained using the Dykstra-like algorithm (Bauschke:2008c, , Theorem 3.3). This algorithm is applied to the proximal subproblem.
$\min_{X\in\mathbbm{R}^{n\times m}}\frac{1}{2}\|X-Z\|^{2}_{F}+{\lambda_{\rm row}}\sum_{i=1}^{n}\|X^{(i)}\|_{2}+{\lambda_{\rm col}}\sum_{j=1}^{m}\|X_{(j)}\|_{2},$ (40)
@@
$\left\{\begin{array}[]{l}Y_{\ell}={\bf prox}_{g_{1}}(Z_{\ell}+P_{\ell})\\ P_{\ell+1}=Z_{\ell}+P_{\ell}-Y_{\ell}\end{array}\right.\quad\mbox{and}\quad\left\{\begin{array}[]{l}Z_{\ell+1}={\bf prox}_{g_{2}}(Y_{\ell}+Q_{\ell})\\ Q_{\ell+1}=Y_{\ell}+Q_{\ell}-Z_{\ell+1},\end{array}\right.$ (41)
where $g_{1}(X)=\lambda_{\rm col}\sum_{j=1}^{m}\|X_{(j)}\|_{2}$ and $g_{2}(X)=\lambda_{\rm row}\sum_{i=1}^{n}\|X^{(i)}\|_{2}$. Recall that:
@@
Using now the definition of $Y_{\ell}$, as presented in (41), we have that $Z_{\ell}+P_{\ell}-Y_{\ell}\in\partial g_{1}(Y_{\ell})$. This observation, combined with Proposition 2, yields
$Z_{\ell}+P_{\ell}-Y_{\ell}\in\partial_{\epsilon_{\ell}}g_{1}(Z_{\ell+1}),\quad\mbox{where}\quad\epsilon_{\ell}=g_{1}(Z_{\ell+1})-g_{1}(Y_{\ell})-\langle{Z_{\ell}+P_{\ell}-Y_{\ell}},{Z_{\ell+1}-Y_{\ell}}\rangle\geq 0.$
On the other hand, it follows from the definition of $Z_{\ell+1}$ in (41) that
$Y_{\ell}+Q_{\ell}-Z_{\ell+1}\in\partial g_{2}(Z_{\ell+1}).$
Thus, it follows from the last two inclusions and the definition of $Z$ that
$V_{\ell}:=Z_{\ell}+P_{\ell}+Q_{\ell}-Z\in\partial_{\epsilon_{\ell}}g(Z_{\ell+1})+Z_{\ell+1}-X_{\ell}+\nabla f(X_{\ell}).$
Therefore, if the Dykstra-like algorithm is stopped when
$g(Z_{\ell}-V_{\ell})-g(Z_{\ell})-\langle\nabla f(X_{\ell}),V_{\ell}\rangle+\frac{(1+\gamma_{1})}{2}\|V_{\ell}\|^{2}+(1+\gamma_{2})\varepsilon_{\ell}\leq\frac{(1-\tau-\alpha)}{2}\|X_{\ell}-Z_{\ell}\|^{2},$
then the triple $(\tilde{X}_{k},V_{k},\varepsilon_{k}):=(Z_{\ell},V_{\ell},\epsilon_{\ell})$ satisfies (10)-(11).
@@
Considering that the IPG-ELS method integrates two effective strategies: (i) permitting the subproblem to be solved inexactly to meet a relative approximation criterion, and (ii) employing an explicit linesearch procedure that computes the proximal operator only once per iteration, our primary goal is to demonstrate that, in certain cases, the proposed method surpasses the proximal gradient method that employs only one of these strategies. Consequently, we compare the new algorithm with two alternative schemes: an exact version of the IPG-ELS method, denoted by PG-ELS, which corresponds to IPG-ELS with $\gamma_{1}=\gamma_{2}=0$, $\theta=0.5$, $\tau=1$, and $\max\{\|V_{k}\|,\epsilon_{k}\}\leq 10^{-12}$, replacing the inexact criterion in (11); and an IPG method with a fixed Stepsize, corresponding to (Millan:2019, , Algorithm 2) with $\alpha_{k}=1/L$ and $\sigma=0.9$, where $L$ is the Lipschitz constant of $f$. This algorithm is denoted by IPG-FixStep and is defined as
$X_{k+1}=\tilde{X}_{k}-V_{k},\quad\forall k\geq 0,$
where $\tilde{X}_{k}$ and $V_{k}$ satisfy
$LV_{k}\in\partial_{\varepsilon_{k}}g(\tilde{X}_{k})+L(\tilde{X}_{k}-X_{k})+\nabla f(X_{k}),\quad\|V_{k}\|^{2}+\frac{2\varepsilon_{k}}{L}\leq 0.9\|X_{k}-\tilde{X}_{k}\|^{2}.$
@@
The initialization parameters for the IPG-ELS method were set as $\tau=0.8$, $\theta=0.5$, $\gamma_{1}=\gamma_{2}=1.1$, and $\alpha=0.01$. For all tests, we initialized $X_{0}=0\in\mathbbm{R}^{n\times m}$, and set $\lambda_{\text{row}}=\lambda_{\text{col}}=0.01$. The IPG-ELS method was executed for $101$ outer iterations to establish a baseline performance metric, $F_{*}:=F(X_{101})$. The other two algorithms were terminated as soon as $F(x_{k})\leq F_{*}$. The algorithms were evaluated on six datasets from Cano:2005 ; Guyon:2004 ; Team:2008 : Colon tumor ($62\times 2000$), heart disease ($303\times 14$), central nervous system (CNS) ($60\times 7129$), lung cancer-Michigan ($96\times 7129$), Secom ($1567\times 590$), and Cina0 ($132\times 16033$).

### v2

@@ 4 Inexact Proximal Gradient Method
In the following, we formally describe the proposed method.
@@
Note that the novel explicit linesearch in Step 4 of the IPG-ELS is related to the ones proposed in Salzo:2017a ; BelloCruz:2016a when $v_{k}=0$ and $\varepsilon_{k}=0$. Moreover, note that the linesearch does not evaluate the inexact proximal subproblem inside the inner loop of Step 4. Hence, only one inexact proximal solution is computed per iteration.
@@
The above lemma ensures that the linesearch procedure in Step 4 of the IPG-ELS method terminates after a finite number of iterations. This result is fundamental to the convergence analysis of the method, as it guarantees that the linesearch procedure is well-defined.
Next, we provide a remark on the possibility of initiating the linesearch with a larger stepsize.
Remark 2 (Over-relaxation Strategy)
In the current implementation, the linesearch algorithm begins with an initial stepsize $\beta=1$. Considering larger initial values for $\beta$, thereby adopting an over-relaxation strategy, may potentially accelerate convergence by extrapolating beyond standard update steps. This approach is related to the concept of conic averagedness BartzDaoPhan , which offers a framework for analyzing the convergence properties of fixed-point algorithms under relaxed conditions. However, initiating the linesearch with $\beta>1$ presents significant challenges to our current approach. While the linesearch remains well-defined for any $\beta>1$, and the proof of finite termination would be directly extended, the convergence and complexity analysis of our scheme depend on the convexity of the function $f$ and the appropriateness of the direction $y_{k}-x_{k}$ to ensure a sufficient decrease in the objective function. Therefore, our convergence analysis does not cover such cases, and the practical implications for this approach are unclear.
The subsequent analysis investigates the complexity of the linesearch procedure introduced in Step 4 of the IPG-ELS method. Here we assume that the gradient of the function $f$, $\nabla f$, is Lipschitz continuous only to establish an upper bound for the number of iterations required by the linesearch procedure in Step 4.
@@
This lemma provides a sufficient condition to ensure that the lower bound of the sequence generated by our linesearch is strictly greater than 00. Specifically, if $\nabla f$ is $L$-Lipschitz continuous, then the stepsizes $\beta_{k}$, produced through the linesearch (34), are guaranteed to be bounded below by a positive constant $\beta>0$, i.e., $\beta_{k}\geq\beta$ for all $k\in\mathbb{N}$. Moreover, it is possible to relax the global Lipschitz condition to something local, such as $\nabla f$ being locally Lipschitz continuous around any solution of problem (1), as was done in Proposition 5.4(ii) of BelloCruz:2016a . In fact, the proof of Lemma 6 may be readily adapted to establish the same complexity now with respect to the locally Lipschitz constant $\mathcal{L}$.
Lemma 7 (Locally Lipschitz Condition and Linesearch Complexity)
If $\nabla f$ is $\mathcal{L}$-locally Lipschitz continuous at any solution of problem (1), then any $\beta\leq\tau/(2\mathcal{L})$ satisfies (34). Consequently, the linesearch procedure in Step 4 of the IPG-ELS method terminates after at most
$\ell:=\left\lceil\frac{\ln(\min\{\tau/(2\mathcal{L}),1\})}{\ln(\theta)}\right\rceil$ (39)
iterations.
Note that the assumption of the gradient of $f$ in problem (1) being locally, rather than globally, Lipschitz continuous is commonly encountered in practice. For instance, this condition arises in the Poisson linear inverse regularization problem with Kullback–Leibler divergence Csiszar:1991 ; Vardi:1985 . We emphasize that this weaker assumption is required exclusively for analyzing the complexity of the linesearch procedure, as it ensures that the stepsizes generated by the proposed linesearch are uniformly bounded away from zero. Additionally, it may be used as sufficient condition for establishing the convergence rate of the functional value sequence. It is worth noting that the finite termination of the linesearch and convergence of the IPG-ELS method do not rely on such assumption.
5 Convergence and Complexity Analysis of the IPG-ELS Method
In this section, we focus on analyzing the convergence properties of the IPG-ELS method. We establish the convergence and its iteration complexity in terms of functional values and a residual associated with an approximate solution, as defined in Definition 1.
@@ 5 Convergence and Complexity Analysis of the IPG-ELS Method
Additionally, the sequence $(F(x_{k}))_{k\in\mathbbm{N}}$ is decreasing and convergent, and $\sum_{k=0}^{+\infty}\|x_{k}-x_{k+1}\|^{2}<+\infty$.
@@
So, the sequence $(F(x_{k}))_{k\in\mathbbm{N}}$ is decreasing and convergent because it is bounded from below by $F(x_{*})$. Moreover, the last inequality implies that $\sum_{k=0}^{+\infty}\|x_{k}-x_{k+1}\|^{2}<+\infty$. ∎
Next, we establish the full convergence of the sequence $(x_{k})_{k\in\mathbbm{N}}$ to a solution of problem (1). The proof is based on the quasi-Fejér convergence of the sequence $(x_{k})_{k\in\mathbbm{N}}$ to the set $S_{*}$, as defined in Definition 2. Note that to establish the convergence of $(x_{k})_{k\in\mathbbm{N}}$ we do not require any local or global Lipschitz continuity assumption of $\nabla f$.
@@
$\sum_{k=0}^{+\infty}\delta_{k}=$ $2\sum_{k=0}^{+\infty}\Big{[}F(x_{k})-F(x_{k+1})\Big{]}\leq 2\Big{[}F(x_{0})-\lim_{k\to+\infty}F(x_{k+1})\Big{]}$
@@
This together with (44) tells us that the sequence $(x_{k})_{k\in\mathbbm{N}}$ is quasi-Fejér convergent to $S_{*}$ via Definition 2. By Lemma 1(a), the sequence $(x_{k})_{k\in\mathbbm{N}}$ is bounded. Let $\bar{x}$ be an accumulation point of $(x_{k})_{k\in\mathbbm{N}}$. Hence, there exists a subsequence $(x_{\ell_{k}})_{k\in\mathbbm{N}}$ converging to $\bar{x}$.
@@
Summing for $k=0,\ldots,m$, the above inequality implies that
@@
$\beta\sum_{k=0}^{+\infty}\left[F(x_{\ell_{k}})-F(x_{*})\right]\leq\sum_{k=0}^{+\infty}\beta_{\ell_{k}}\left[F(x_{\ell_{k}})-F(x_{*})\right]\leq\sum_{k=0}^{+\infty}\beta_{k}\left[F(x_{k})-F(x_{*})\right]<+\infty,$
@@
$F(x_{*})\leq F(\bar{x})\leq\liminf_{k\rightarrow+\infty}F(x_{\ell_{k}})=\lim_{k\rightarrow+\infty}F(x_{\ell_{k}})=F(x_{*}),$
@@
therefore, $\lim_{k\rightarrow+\infty}\,\|x_{\ell_{k}}-\tilde{x}_{\ell_{k}}\|=0$ and $\lim_{k\rightarrow+\infty}\|v_{\ell_{k}}\|=0$. Additionally, we can use (35) to show that $\lim_{k\rightarrow+\infty}\varepsilon_{\ell_{k}}=0$. Thus, $\bar{x}$ is also an accumulation point of the sequence $(\tilde{x}_{k})_{k\in\mathbb{N}}$, and $\tilde{x}_{\ell_{k}}\to\bar{x}$ as $k\to+\infty$. Moreover, we have that
@@
We continue by showing the convergence rate of the functional values sequence $(F(x_{k}))_{k\in\mathbbm{N}}$.
Theorem 5.2 (Convergence Rate of the IPG-ELS Method)
Summing the above inequality over $\ell=0,1,\ldots,k$, we have
$(k+1)\left(F(x_{k})-F_{*}\right)\leq\frac{1}{2\beta}\left(\|x_{0}-x_{*}\|^{2}+2\left[F(x_{0})-F_{*}\right]\right).$
Since $x_{*}\in S_{*}$ is arbitrary, the proof of (51) follows. ∎
Remark 3 (Complexity of $\eta$-Approximate Solution)
It follows from Theorem 5.2 that, given any $\eta>0$, the IPG-ELS method generates an $\eta$-approximate solution $x_{k}$ to problem (1), in the sense that $F(x_{k})-F_{*}\leq\eta$ in at most $k=\mathcal{O}(1/\eta)$ iterations. We note further that Lemma 7 guarantees that the locally Lipschitz assumption for the gradient of $f$ can be used as a sufficient condition for establishing the convergence rate in Theorem 5.2.
We end this section by proving the complexity of an $\eta$-approximate stationary solution for problem (1) as in Definition 1. For this complexity result, we assume that $\nabla f$ is $L$-Lipschitz continuous on the domain of $g$.
Theorem 5.3 (Complexity of $\eta$-Approximate Stationary Solution)
Additionally, if $\alpha\in(0,1-\tau)$, and $\nabla f$ is $L$-Lipschitz continuous on ${\bf dom\,}g$, then, given a tolerance $\eta>0$, the IPG-ELS method generates an $\eta$-approximate stationary solution $\tilde{x}_{k}$ to problem (1) with residues $(w_{k},\varepsilon_{k})$, in the sense of Definition 1, in at most $k=O\left(1/\eta^{2}\right)$ iterations.
$\|x_{{\ell_{k}}}-\tilde{x}_{{\ell_{k}}}\|^{2}\leq\frac{d_{0}^{2}+2[F(x_{0})-F(x_{*})]}{\alpha\beta(k+1)}.$ (55)
Now, from the definition of $w_{\ell}$, (56), the Cauchy-Schwarz inequality, the fact that $x_{k},\tilde{x}_{k}$ are in the domain of $g$, and the $L$-Lipschitz continuity of $\nabla f$ on the domain of $g$, we have, for every $\ell\in\mathbbm{N}$,
Moreover, it follows from (56) that
which in turn implies that
@@ 6 Numerical Experiments
In this section, we investigate the numerical behavior of the IPG-ELS method in solving the CUR-like factorization optimization problem JMLR:v12:mairal11a . Consider $\mathbbm{E}=\mathbbm{R}^{n\times m}$. Given a matrix $W\in\mathbbm{R}^{m\times n}$, the objective is to find a matrix with sparse rows and columns, $X\in\mathbbm{R}^{n\times m}$, such that $WXW$ approximates $W$. This problem can be formulated as the following splitting optimization problem:
$\min_{X\in\mathbbm{R}^{n\times m}}\left\{\frac{1}{2}\|W-WXW\|_{F}^{2}+\lambda_{\text{row}}\sum_{i=1}^{n}\|X^{(i)}\|_{2}+\lambda_{\text{col}}\sum_{j=1}^{m}\|X_{(j)}\|_{2}\right\},$ (60)
@@
In this case, the gradient of $f$ is given by $\nabla f(X)=W^{T}(WXW-W)W^{T}$ and has a Lipschitz constant $L=\|W^{T}W\|_{F}^{2}$. Note that the proximal operator of $g$ does not have a closed-form solution; however, $g$ exhibits a separable structure. By following the approach described in Section 3.1, we compute a triple $(\tilde{X}_{k},0,\varepsilon_{k})$ that satisfies conditions (32)-(33) using the Dykstra-like algorithm (Bauschke:2008c, , Theorem 3.3). This algorithm is applied to the proximal subproblem
$\min_{X\in\mathbbm{R}^{n\times m}}\left\{\frac{1}{2}\|X-Z\|^{2}_{F}+{\lambda_{\rm row}}\sum_{i=1}^{n}\|X^{(i)}\|_{2}+{\lambda_{\rm col}}\sum_{j=1}^{m}\|X_{(j)}\|_{2}\right\},$ (61)
@@
$\left\{\begin{array}[]{l}y_{\ell}={\bf prox}_{g_{1}}(z_{\ell}+p_{\ell})\\ p_{\ell+1}=z_{\ell}+p_{\ell}-y_{\ell}\end{array}\right.\quad\mbox{and}\quad\left\{\begin{array}[]{l}z_{\ell+1}={\bf prox}_{g_{2}}(y_{\ell}+q_{\ell})\\ q_{\ell+1}=y_{\ell}+q_{\ell}-z_{\ell+1},\end{array}\right.$
where $g_{1}(x)=\lambda_{\rm col}\sum_{j=1}^{m}\|x_{(j)}\|_{2}$ and $g_{2}(x)=\lambda_{\rm row}\sum_{i=1}^{n}\|x^{(i)}\|_{2}$. It follows from the separable property of the proximal operator and Section 6.5.1 of Parikh-Boyd:2014 that
@@
Hence, if the Dykstra-like algorithm is terminated when
$\varepsilon_{\ell}\leq\frac{(1+\gamma_{2})(1-\tau-\alpha)}{2}\|X_{k}-z_{\ell}\|^{2},$
then $\tilde{X}_{k}:=z_{\ell}$ is an inexact proximal solution of (61) with residual $(V_{k},\varepsilon_{k}):=(0,\varepsilon_{\ell})$, where $\varepsilon_{\ell}$ is as defined in (17).
@@
Considering that the IPG-ELS method integrates two effective strategies: (i) permitting the subproblem to be solved inexactly to meet a relative approximation criterion, and (ii) employing an explicit linesearch procedure that computes the proximal operator only once per iteration, our primary goal is to demonstrate that, in certain cases, the proposed method surpasses the proximal gradient method that employs only one of these strategies. Consequently, we compare the new algorithm with three alternative schemes: an exact version of the IPG-ELS method, denoted by PG-ELS, which corresponds to IPG-ELS with $\gamma_{1}=\gamma_{2}=0$, $\theta=0.5$, $\tau=1$, and $\epsilon_{k}\ \leq 10^{-12}$, replacing the inexact criterion in (33) with $v_{k}=0$; an IPG method with a fixed Stepsize, corresponding to (Millan:2019, , Algorithm 2) with $\alpha_{k}=1/L$, $\bar{w}_{k}=L(X_{k}-\tilde{X}_{k})-\nabla f(X_{k})$ and $\sigma^{2}=0.9$, where $L$ is the Lipschitz constant of $f$. This algorithm is denoted by IPG-FixStep and is defined as $X_{k+1}=\tilde{X}_{k},$ for all $k\geq 0,$ where $\tilde{X}_{k}$ satisfies
$0\in\partial_{\varepsilon_{k}}g(\tilde{X}_{k})+L(\tilde{X}_{k}-X_{k})+\nabla f(X_{k}),\quad{\varepsilon_{k}}\leq{0.9L}\|\tilde{X}_{k}-\left(X_{k}-(1/L)\nabla f(x_{k})\right)\|^{2}/2;$
and an instance of Tseng’s modified forward-backward splitting method, as described in Monteiro:2010 , applied to (60). The stepsize is fixed at $0.9/L$, and the triple $(\tilde{X}_{k},0,\varepsilon_{k})$, associated with the proximal subproblem, is computed to ensure $\varepsilon_{k}\leq 10^{-12}$. This algorithm will be referred to as Tseng-MFBS.
@@
For all methods, the approximate proximal solution $\tilde{X}_{k}$ of (61), along with its residual $(0,\varepsilon_{k})$, is computed using the Dykstra-like algorithm as described above. The initialization parameters for the IPG-ELS method were set as $\tau=0.8$, $\theta=0.5$, $\gamma_{1}=\gamma_{2}=1.1$, and $\alpha=0.01$. For all tests, we initialized $X_{0}=0\in\mathbbm{R}^{n\times m}$, and set $\lambda_{\text{row}}=\lambda_{\text{col}}=0.01$. The IPG-ELS method was executed for $101$ outer iterations to establish a baseline performance metric, $F_{*}:=F(X_{101})$. The other three algorithms were terminated as soon as $F(X_{k})\leq F_{*}$ or after reaching a maximum of 2001 iterations.. The algorithms were evaluated on six datasets from Cano:2005 ; Guyon:2004 ; Team:2008 : Colon tumor ($62\times 2000$), heart disease ($303\times 14$), central nervous system (CNS) ($60\times 7129$), lung cancer-Michigan ($96\times 7129$), Secom ($1567\times 590$), and Cina0 ($132\times 16033$).